OpenStudy (anonymous):

how many qt/min are in 9 gal/s?

OpenStudy (wolf1728):

How many quarts (is that minims?) in 9 gallons ? waht's the /s at the very end?

OpenStudy (anonymous):

I have no clue that's why I asked

OpenStudy (anonymous):

its a rate question, probably more like, what is the rate of 9gal/sec in terms of qt per minute

OpenStudy (anonymous):

\[\frac{ 9gal }{ \sec}\times \frac{ 60\sec }{ 1 \min}\times \frac{ 1 qt }{ 0.25 gal} = \frac{ ?qt }{ \min }\]

OpenStudy (anonymous):

2160 qt/min

OpenStudy (wolf1728):

wow - and I thought they were asking for minims. (one minim is 1/480 of a fluid ounce) Okay - well good work folks.

OpenStudy (wolf1728):

Wow - 9 gallons per second? That's a HUGE volume.

OpenStudy (wolf1728):

That's 1.2031cubic feet per second and a cubic foot is 62.4 pounds. So, that is 75 POUNDS of water a second That is a TON of water every 27 seconds!!! I guess that's not coming from a kitchen faucet!